OpenStudy (yinshy):

What value of m makes the statement true? –(m + 15) = –18 + (–15) A. –33 B. –18 C. 3 D. 18

OpenStudy (yinshy):

@Awolflover1

OpenStudy (shaima):

its 18 lmao

Awolflover1 (awolflover1):

\[−m−15+15=−33+15\]

Awolflover1 (awolflover1):

Add 15 to both sides

Awolflover1 (awolflover1):

@shaima Incorrect

OpenStudy (yinshy):

sorry i keep not reading my question befor i post it

OpenStudy (yinshy):

A

Awolflover1 (awolflover1):

Nope Not A

OpenStudy (yinshy):

hmmmm choose wizly

OpenStudy (yinshy):

bc bc bc bc bc i say b

Awolflover1 (awolflover1):

\[-33+15=?\]

OpenStudy (yinshy):

-18

Awolflover1 (awolflover1):

yep there we go :)

OpenStudy (shaima):

nah its 18

OpenStudy (yinshy):

bruh a master vs a student what are the ods of the master being wong???

Awolflover1 (awolflover1):

rightt >.< \[−m −1 = −18 −1\]

Awolflover1 (awolflover1):

m=18 oops xD

OpenStudy (yinshy):

now im lost

OpenStudy (yinshy):

so she was right

OpenStudy (yinshy):

>_< i trused you sensei how could you do this to me

Awolflover1 (awolflover1):

yes but remember no direct answers lol @shaima cuz that was confusing

OpenStudy (yinshy):

JK

OpenStudy (shaima):

lol i didnt my math i was like howw am i wrong

OpenStudy (shaima):

did*

OpenStudy (yinshy):

we are soz

Awolflover1 (awolflover1):

but you're suppose to explain on here LOL

OpenStudy (yinshy):

ok now last question